# Tokenization Is More Than Compression

### v1

@@ 6.2 Vocabulary Construction
One way to examine the effects of vocabulary construction is to compare the resulting vocabularies of top-down methods trained using an initial vocabulary to the method itself. Figure 5 presents an area-proportional Venn diagram of the overlap in 40,960-sized vocabularies between BPE (6) and variants of PathPieceL (1) and SaGe (6) that were trained using an initial BPE vocabulary of size $1442^{18}=262,$ 144.131313See Figure 10 in Appendix E.2 for analogous results for Unigram, which behaves similarly. While BPE and PathPieceL overlap considerably, SaGe produces a more distinct set of tokens.
@@
The difference in tie breaking strategy in PathPiece between longest-token (L) and random tiebreaks (R) turns out not to be significant (see Appendix E). Further discussion of the effects of vocabulary construction through a random baseline is provided in Appendix F.
@@ 6.3 Initial Vocabulary
PathPiece, SaGe, and Unigram all require an initial vocabulary. [^14] The HuggingFace Unigram implementation starts with the one million most top $n$-grams, but sorted according to the count times the length of the token, introducing a bias toward longer tokens. For PathPiece and SaGe, we experimented with initial vocabularies of size 262,144 constructed from either the most frequent $n$-grams, or trained using either BPE or Unigram. For PathPieceL, using a BPE initial vocabulary (1) is statistically better than both Unigram (9) and $n$-grams (16), with $p\leq 0.01$. Using an $n$-gram initial vocabulary leads to the lowest performance, with statistical significance. Comparing ranks 6, 8, and 10 reveals the same pattern for SaGe, although the difference between 8 and 10 is not significant.
6.4 Segmentation
@@ 7 Conclusion
We investigated the hypothesis that reducing the corpus token count (CTC) would improve downstream performance. We find that five different tokenizers, all with different CTCs, can perform comparably (§5.2). Moreover, when comparing CTC and downstream accuracy across all our experimental settings in Figure 3, we do not find a clear relationship between the two.
This casts doubt on the current understanding of why BPE is particularly effective. Our findings dovetail with Bostrom and Durrett (2020)’s results that unigram tokenizers better align with morphological segmentation compared to BPE tokenizers, further suggesting there is not a straightforward explanation of what makes a tokenizer effective.
Finally, this work contributes a number of practical contributions: (1) BPE initial vocabularies work best for top-down vocabulary constructions; (2) Space pre-tokenization Gow-Smith et al. (2022) outperforms the more common FirstSpace pre-tokenization for PathPiece; (3) overall, vocabulary size has little impact on downstream performance. To further encourage research in this direction, we make all of our tokenized vocabularies publicly available, along with the model weights from our 64 language models.
@@ Acknowledgments
Thanks to Charles Lovering at Kensho for his insightful suggestions, and to Michael Krumdick, Mike Arov, and Brian Chen at Kensho for their help with the language model development process. This research was supported in part by the Israel Science Foundation (grant No. 1166/23).
@@ Appendix C Description of Downstream Tasks
Appendix D Effect of model size and model convergence
Figure 6: 40,960 vocab average accuracy at various models sizes
@@ Appendix E Additional Analysis
E.1 Digit Pre-tokenization
@@ Appendix F RandTrain
In line with Subsection 6.4, Greedy performs significantly better than PathPieceL segmentation in all 3 cases. However, for the two cases with an $n$-gram initial vocabulary the PathPieceL segmentation did extremely poorly. The RandTrain vocabulary construction, $n$-gram initial vocabulary, and PathPieceL segmentation interact somehow to give accuracies well below any others.
@@ Appendix G Detailed Experimental Results
This section gives the detailed accuracy results for the 10 downstream evaluation tasks on each model that was trained. The tables are divided by the vocabulary size used, with Table 3 and Table 4 for 32,768; Table 5 and Table 6 for 40,960; and Table 7 and Table 8 for 49,152. The highest value or values (in the case of ties) are shown in bold. Table 9 show the same results as Table 1, but are sorted from best to worst by rank.
@@
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
@@
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
@@
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
@@
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273

### v2

@@ 6.2 Vocabulary Construction
One way to examine the effects of vocabulary construction is to compare the resulting vocabularies of top-down methods trained using an initial vocabulary to the method itself. Figure 5 presents an area-proportional Venn diagram of the overlap in 40,960-sized vocabularies between BPE (6) and variants of PathPieceL (1) and SaGe (6) that were trained using an initial BPE vocabulary of size $1442^{18}=262,$ 144.161616See Figure 12 in Appendix E.3 for analogous results for Unigram, which behaves similarly. While BPE and PathPieceL overlap considerably, SaGe produces a more distinct set of tokens.
@@ 6.3 Initial Vocabulary
PathPiece, SaGe, and Unigram all require an initial vocabulary. [^17] The HuggingFace Unigram implementation starts with the one millionp $n$-grams, but sorted according to the count times the length of the token, introducing a bias toward longer tokens. For PathPiece and SaGe, we experimented with initial vocabularies of size 262,144 constructed from either the most frequent $n$-grams, or trained using either BPE or Unigram. For PathPieceL, using a BPE initial vocabulary (1) is statistically better than both Unigram (9) and $n$-grams (16), with $p\leq 0.01$. Using an $n$-gram initial vocabulary leads to the lowest performance, with statistical significance. Comparing ranks 6, 8, and 10 reveals the same pattern for SaGe, although the difference between 8 and 10 is not significant.
6.4 Effect of Model Size
Figure 6: 40,960 vocab average accuracy at various models sizes
@@ 7 Conclusion
We investigate the hypothesis that reducing the corpus token count (CTC) would improve downstream performance, as suggested by Gallé (2019) and Goldman et al. (2024) when they varied aspects of BPE. When comparing CTC and downstream accuracy across all our experimental settings in Figure 3, we do not find a clear relationship between the two. We expand on the findings of Ali et al. (2024) who did not find a strong relation when comparing 3 tokenizers, as we run 18 experiments varying the tokenizer, initial vocabulary, pre-tokenizer, and inference method. Our results suggest compression is not a straightforward explanation of what makes a tokenizer effective.
Finally, this work makes several practical contributions: (1) vocabulary size has little impact on downstream performance over the range of sizes we examined (§5.1); (2) five different tokenizers all perform comparably, with none outperforming at statistical significance (§5.2); (3) BPE initial vocabularies work best for top-down vocabulary construction (§6.3). To further encourage research in this direction, we make all of our trained vocabularies publicly available, along with the model weights from our 64 language models.
@@ Acknowledgments
Thanks to Charles Lovering at Kensho for his insightful suggestions, and to Michael Krumdick, Mike Arov, and Brian Chen at Kensho for their help with the language model development process. This research was supported in part by the Israel Science Foundation (grant No. 1166/23). Thanks to an anonymous reviewer who pointed out the large change in CTC when comparing Huggingface BPE and Unigram, in contrast to the previous literature using the SentencePiece implementations Kudo and Richardson (2018).
@@ Appendix C Description of Downstream Tasks
Appendix D Effect of model convergence
@@ Appendix E Additional Analysis
E.1 Segmentation
Figure 8: Segmentation of BPE.
Pairwise $p$-values between the pairs of runs are
$p$(3,4)=0.52, $p$(3,13)=4.4e-5, $p$(4,13)=8.8e-6.
Figure 9: Segmentation of Unigram.
Pairwise $p$-values between the pairs of runs are
$p$(2,7)=0.041, $p$(2,17)=2.9e-06, $p$(7,17)=2.9e-06
E.2 Digit Pre-tokenization
@@ Appendix F RandTrain
In line with Subsection E.1, Greedy performs significantly better than PathPieceL segmentation in all 3 cases. However, for the two cases with an $n$-gram initial vocabulary the PathPieceL segmentation did extremely poorly. The RandTrain vocabulary construction, $n$-gram initial vocabulary, and PathPieceL segmentation interact somehow to give accuracies well below any others.
@@ Appendix G Detailed Experimental Results
This section gives the detailed accuracy results for the 10 downstream evaluation tasks on each model that was trained. The tables are divided by the vocabulary size used, with Table 4 and Table 5 for 32,768; Table 6 and Table 7 for 40,960; and Table 8 and Table 9 for 49,152. The highest value or values (in the case of ties) are shown in bold. Table 10 show the same results as Table 1, but are sorted from best to worst by rank. The corpus token count (CTC), Rényi efficiencies, and average accuracies for the 54 runs in Figure 3 are given in Table 11.
@@
Vocab Constr Init Voc Pre-tok Segment Avg arc_easy copa mktg mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segment qa4mre race sciq sociology wsc273
@@
Vocab Constr Init Voc Pre-tok Segment Avg arc_easy copa mktg mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segment qa4mre race sciq sociology wsc273
@@
Vocab Constr Init Voc Pre-tok Segment Avg arc_easy copa mktg mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segment qa4mre race sciq sociology wsc273
@@
Rank Vocab Size Avg Acc CTC Eff $\alpha$=1.5 Eff $\alpha$=2 Eff $\alpha$=2.5 Eff $\alpha$=3 Eff $\alpha$=3.5
1 32,768 49.3 1.48 0.604 0.516 0.469 0.441 0.422
1 40,960 49.4 1.46 0.589 0.503 0.457 0.429 0.411
1 49,152 49.4 1.44 0.578 0.492 0.448 0.420 0.402
2 32,768 49.2 1.79 0.461 0.371 0.324 0.295 0.277
2 40,960 49.1 1.77 0.451 0.362 0.316 0.289 0.271
2 49,152 48.8 1.76 0.444 0.356 0.311 0.284 0.266
3 32,768 48.8 1.52 0.594 0.505 0.459 0.431 0.414
3 40,960 50.0 1.49 0.579 0.491 0.446 0.420 0.403
3 49,152 48.1 1.47 0.567 0.481 0.437 0.411 0.394
4 32,768 48.3 1.50 0.605 0.517 0.471 0.442 0.423
4 40,960 49.1 1.48 0.590 0.504 0.458 0.430 0.412
4 49,152 49.5 1.46 0.579 0.494 0.449 0.421 0.403
5 32,768 48.5 1.54 0.598 0.507 0.461 0.433 0.415
5 40,960 49.1 1.51 0.583 0.494 0.448 0.421 0.404
5 49,152 48.8 1.49 0.571 0.483 0.439 0.412 0.396
6 32,768 47.9 1.78 0.545 0.466 0.422 0.396 0.378
6 40,960 49.2 1.76 0.533 0.455 0.413 0.387 0.369
6 49,152 48.7 1.75 0.523 0.447 0.405 0.379 0.362
7 32,768 47.9 1.81 0.510 0.431 0.387 0.359 0.340
7 40,960 48.5 1.79 0.500 0.423 0.381 0.354 0.335
7 49,152 48.6 1.77 0.493 0.416 0.375 0.348 0.330
8 32,768 47.5 1.63 0.629 0.536 0.482 0.447 0.424
8 40,960 48.5 1.62 0.615 0.524 0.470 0.437 0.415
8 49,152 48.0 1.62 0.605 0.515 0.462 0.429 0.407
9 32,768 46.9 1.74 0.508 0.419 0.372 0.343 0.323
9 40,960 48.5 1.72 0.491 0.403 0.356 0.328 0.309
9 49,152 48.4 1.72 0.477 0.389 0.343 0.315 0.296
10 32,768 48.4 2.02 0.485 0.409 0.366 0.339 0.320
10 40,960 46.9 2.01 0.474 0.401 0.358 0.331 0.313
10 49,152 47.8 2.01 0.466 0.393 0.352 0.325 0.307
11 32,768 48.4 1.77 0.587 0.512 0.470 0.443 0.425
11 40,960 46.9 1.76 0.575 0.501 0.460 0.433 0.415
11 49,152 47.2 1.76 0.565 0.492 0.452 0.426 0.408
12 32,768 47.5 2.33 0.236 0.164 0.138 0.124 0.116
12 40,960 45.4 2.30 0.228 0.159 0.133 0.120 0.112
12 49,152 47.3 2.29 0.223 0.155 0.130 0.117 0.109
13 32,768 45.6 1.50 0.606 0.518 0.470 0.442 0.423
13 40,960 46.7 1.47 0.591 0.504 0.458 0.430 0.412
13 49,152 47.2 1.45 0.579 0.494 0.449 0.421 0.403
14 32,768 45.3 1.46 0.616 0.532 0.490 0.465 0.448
14 40,960 45.8 1.43 0.602 0.519 0.478 0.453 0.437
14 49,152 45.5 1.42 0.591 0.508 0.468 0.444 0.428
15 32,768 44.6 1.47 0.620 0.533 0.490 0.464 0.447
15 40,960 44.9 1.44 0.605 0.520 0.478 0.453 0.436
15 49,152 45.0 1.42 0.594 0.509 0.468 0.443 0.427
16 32,768 44.8 1.36 0.677 0.571 0.514 0.480 0.457
16 40,960 45.5 1.33 0.662 0.556 0.500 0.466 0.444
16 49,152 43.9 1.31 0.650 0.544 0.489 0.456 0.435
17 32,768 43.6 1.77 0.471 0.380 0.333 0.304 0.285
17 40,960 43.1 1.75 0.462 0.372 0.326 0.298 0.280
17 49,152 44.0 1.74 0.455 0.366 0.320 0.293 0.275
18 32,768 43.5 1.29 0.747 0.617 0.549 0.511 0.486
18 40,960 44.0 1.26 0.736 0.603 0.535 0.497 0.474
18 49,152 42.2 1.25 0.728 0.591 0.524 0.487 0.464
Table 11: Average Accuracy (%) vs. Corpus Token Count (CTC, in billions) by vocabulary size, for Figure 3. Also includes the corresponding Rényi efficiency (Zouhar et al., 2023a) for various orders $\alpha$.
Vocab Constr Init Voc Pre-tok Segment Avg arc_easy copa mktg mathqa piqa
@@
Vocab Constr Init Voc Pre-tok Segment qa4mre race sciq sociology wsc273